OpenStudy (anonymous):

The measurement of one angle of a right triangle is 42°. What is the measurement of the missing angle?

OpenStudy (simplymarie_x):

It'd be x + 42 + 90 = 180. Do the same as before

OpenStudy (simplymarie_x):

(the 90 comes from it being a right triangles. right angle = 90 degrees)

OpenStudy (anonymous):

so 42 + 90 = 132

OpenStudy (simplymarie_x):

Yes, then do x + 132 = 180 or 180 - 132 = your missing angle

OpenStudy (anonymous):

so x=48

OpenStudy (simplymarie_x):

Good! Yes!